OpenStudy (asapbleh):

Show that the equation x^3 + x^2 -2x = 1 has at least one solution in the interval [-1,1].

OpenStudy (asapbleh):

This is how I did it. But i wonder if it is correct. x^3+x^2-2x=1 x^3-x^2-2x-1=0 x(x^2+x-2-1/x = 0

OpenStudy (anonymous):

Plug in both -1 and 1 for x. If you get one positive and one negative value then the intermediate value theorem states there has to be a value that is 0 somewhere between the two

OpenStudy (asapbleh):

Oic, let me try plugging in.

OpenStudy (asapbleh):

I dont get it.

OpenStudy (anonymous):

was there a y or an f(x) anywhere in the original problem

OpenStudy (asapbleh):

No, the whole question was what I stated in the question,

OpenStudy (anonymous):

Oh then you have to set the expression to 0 and replace 0 with y to get a function. Then plug in -1 and 1 for x and if the values for y are positive and negative tthen the intermediate value theorem proves it. So just take x^3+x^2-2x-1 and plug in -1 and 1

OpenStudy (asapbleh):

Oh, ok. lemme try that.

OpenStudy (asapbleh):

I got -1=0 and 2=0

OpenStudy (anonymous):

but the 0 is a y now so its -1=y and 2=y

OpenStudy (anonymous):

 basically you get this and your function goes between these two points (-1,2) and (1,-1) so it has to pass through the x axis somwhere between -1 and 1

OpenStudy (asapbleh):

Oh, Ic ic, Thanks, I get it now.